A cart with a spring plunger runs into a fixed bar Please include an explanation
Answer #1


we know that change in kinetic energy = k2 -k1 = 1/2 m(v2^2 - v1^2)


from the given data

   A . k2-k1 = 1/2*10(0-4^2) = - 80 J

   B . k2-k1 = 1/2*10((-1^2)-(3^2)) = - 40 J

  
    C. k2-k1 = 1/2*20((-1^2)-(1^2)) = 0 J

   D. k2-k1 = 1/2*5((-3^2)-(5^2)) = - 40 J


so the rank of change in kinetic energies are A >B=D>C